Solve the system of nonlinear equations using substitution or elimination {x^2+y^2=13 (2x-3y=0
Answer #1

hiven: 12 + y2 = 13 0 & 22-3y =0 @ 34 ; ④ = 2 = 2 (39)? +y = 13 १५२ +y?=13 -) 13 y? = 4(13) 4 = y2 =4 y = -2 x=3 y=2 = 2x-312